You wish to prepare a solution of ammonium persulfate (NHJ)S,O) a)Calculate the mass of (NH),5,0, in grams, required to prepare 125.00 mL of 1.50M solution. b) 63.54 mL of a (NHu),S,0, solution of 0.500 M is mixed with 125.00 mL of a (NHJ,s,0, soluti of 1.50 M. Find the concentration of (NH),S,0, in the resulting solution.
Answer #1

a) Molar mass of ( NH4)2S2O8 = 228 g/ mol.

Molarity = W2*1000/M2*V. (1)

Where, W2 = mass of solute[ ( NH4)2S2O8]

M2 = molar of solute = 228 g/mol

V = Volume of solution in mL = 125 mL

Now , putting the above value in the Eq. 1

1.5 = W2*1000/228*125

Or, W2 = 1.5*228*125/1000 = 42.75 g.

Hence, 42.75 g of (NH4)2S2O8 is required.

b) Let final Concentration of the resulting solution after mixing = S3

Applying , the equation

V1S1 + V2S2 = V3S3

63.54*0.500 + 125*1.5 = (125+63.54)*S3

Or, S3 = (31.77 + 187.5)/188.54 = 1.16 M